OpenStudy (anonymous):

Prove that Sin^2(x)tan(x)+cos^2(x)cot(x)+2sin(x)cos(x)=(sin(x)+cos(x))/sin(x)cos(x)

OpenStudy (vishweshshrimali5):

\[\large{\sin^2 x \tan x + \cos^2 x \cot x + 2\sin x \cos x = \cfrac{(\sin x + \cos x)}{\sin x \cos x}}\] Okay this is the question right ?

OpenStudy (anonymous):

Yes

OpenStudy (aum):

Write tan(x) and cot(x) in terms of sin(x) and cos(x). Find the common denominator for the LHS. Add the terms and simplify.

OpenStudy (vishweshshrimali5):

First of all: \[\large{\tan x = \cfrac{\sin x}{\cos x}}\] \[\large{\cot x = \cfrac{1}{\tan x} = \cfrac{\cos x}{\sin x}}\] Use these in the LHS of the identity to be proved

OpenStudy (anonymous):

Can you go a bit further

OpenStudy (kainui):

@Swqaedf You should attempt to use this information he has so helpfully given to see how far you can get! Good luck!

OpenStudy (kirbykirby):

I tried this but I end up with LHS being \[ \frac{1}{\sin x \cos x}\] I tried showing .. which seems to agree with Wolfram which simplified it to \(\csc x \sec x\): http://www.wolframalpha.com/input/?i=sin%5E2x+tan+x+%2B+cos%5E2+x+cot+x+%2B+2+sin+x+cos+x

OpenStudy (vishweshshrimali5):

Okay let me tell you the first few steps: \[\large{\sin^2 (x) \tan (x)+\cos^2 (x) \cot (x)+2\sin (x) \cos (x)}\] \[\large{=\sin^2 x \cfrac{\sin x}{\cos x} + \cos^2 x \cfrac{\cos x}{\sin x} + 2\sin x \cos x}\] \[\large{=\cfrac{\sin^3 x}{\cos x} + \cfrac{\cos^3 x}{\sin x} + 2\sin x \cos x}\] \[\large{= \cfrac{\sin^4 x + \cos^4x + 2\sin^2 x \cos^2 x}{\sin x \cos x}}\] Now we use: \[\large{a^4+b^4+2a^2b^2 = (a^2+b^2)^2}\] Thus the LHS will become: \[\large{=\cfrac{(\sin^2 x + \cos^2 x)^2}{\sin x \cos x}}\] Now you should know that: \[\large{\sin^2 x+ \cos^2 x = 1}\] Use this to simplify the LHS and I tend to agree with @kirbykirby . There appears to be a mistake in the question. The RHS should be: \[\large{\cfrac{1}{\sin x \cos x}}\]